4. Diane has a loan with an interest rate of 6%. The amount of interest is P43.75 for a period of 75 days. How much is her loan using ordinary interest method?



5. Angela read an advertisement for a P50,000 loan. The ad informs the borrower that the company only charges 18% annual interest. If she plans to borrow that amount and repays it at the end of the year, how much is the interest? What is the interest of each month?

1
Expert's answer
2021-11-16T07:38:01-0500

4.

According to ordinary interest method ,

"I = P.r \times (\frac{D}{360})"

I - Interest is the amount paid on the borrowed money by the debtor. (43.75) 

P - Principal is the actual amount taken from the bank/creditor.

r - rate of interest "( 6\% = \frac{6}{100})"

D - Maturity period.(75 days or 0.208 year)


"I = \frac{P\times r\times n }{100}\\

43.75 =( P \times 6\times 0.208) /100\\

P = \frac{(100\times43.75)}{(6\times 0.208)}\\

P = \frac{4375 }{ 1.25} \\

P =3500"

 

The loan amount is 3500.

5.

Interest = Principal × Interest rate

"= 50,000(0.18)\\

= 9000"

Interest per annum = 9000

 

Step 2

Interest per month = Interest per annum / No of months

"= \frac{9000}{12}\\

= 750 \space per\space month"

Interest per month = 750